Let the width(w) and the area(A) be real numbers. The perimeter remains a constant 24cm. Draw a graph for 1<=w<=11 that represents the relationship between the width and the area. Label the axes appropriately
Answer by math_tutor2020(3816)


x = width
P = perimeter
P = 2*length + 2*width
24 = 2*length + 2*width
12 = length + width
12 = length + x
length = 12-x

A(x) = area
A(x) = length*width
A(x) = (12-x)*x
A(x) = -x^2+12x

Graph of A(x) when 1 <= x <= 11

x = width
A(x) = area of the rectangle

Table of selected values
xA(x)
111
220
327
432
535
636
735
832
927
1020
1111

(6,36) is the vertex since it's the highest point.
This vertex point tells us that the area is maxed out at 36 square cm when the width is x = 6 cm.